Circular No. 7171

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1986 U 10 E. Karkoschka, Lunar and Planetary Laboratory, University of Ari-

zona, reports the discovery of a previously unknown satellite of Uranus in seven Voyager 2 images, at a signal that is between 3 and 15 times the background noise. Observed over the 5-day arc 1986 Jan. 18–23, S/1986 U 10 was at apparent mag ∼ 6.5–9.5 (brightening as the spacecraft approached Uranus), corresponding to a mean opposition magnitude of V ∼ 23.6 and diameter roughly 40 km (assuming an albedo similar to that of nearby satellites). With orbital motion 564^{o}.25 ± 0^{o}.10 day^{−1}, an assumed circular orbit in the plane of Uranus’ equator yields an orbital radius of 76 416 km. Conjunction in orbital longitude between S/1986 U 10 and Uranus XIV (Belinda) occurred on 1986 Jan. 24.7 UT.

COMET C/1999 J2 (SKIFF) Additional astrometry and the following very preliminary parabolic

orbital elements appear on MPEC 1999-K08:

T = 1999 Oct. 23.960 TT  ω = 114.^{o}701
Ω = 47.826 }2000.0
q = 7.48613 AU  i = 86.940
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 12 17^{h}00.^{m}39 +63^{o}43.^{′}4 7.390 7.558 95.^{o}7 7.^{o}6 15.1
22 16 46.52 +63 34.8 7.384 7.549 95.5 7.7 15.1
June 1 16 32.51 +63 08.2 7.387 7.541 94.9 7.7 15.1
11 16 19.08 +62 23.3 7.398 7.534 93.8 7.7 15.1
21 16 06.89 +61 21.3 7.416 7.527 92.4 7.8 15.1

V1333 AQUILAE C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence

(OHP), write: “CCD images of V1333 Aql = Aquila X-1 taken with the 1.2-m OHP telescope on May 13.1, 14.1, and 15.1 UT show a rapid increase in mean brightness over this time interval, from V = 17.9 to 17.2. These levels are comparable to the maximum values that we observed during the mid-1996 mini-outburst and the mid-1997 moderate outburst. However, quick-look data from the XTE All-Sky Monitor for May 14–17 show low count rates (< 70 milliCrab). This low x-ray-to-optical luminosity ratio may perhaps signal a repeat of the 1996 mini-outburst.”

1999 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green